Jamie wants to make a 20 lbs mix worth $22, out of $20 per lbs of Kona coffee with a $28 per lbs of Mountain coffee. How much of each mix is needed to make the 20lbs for $22 per lbs?
Answer by rwm(914)

k+m=20
20*22=20*k+28*m